# Advancing Galaxy Analysis: AI-Powered Detection and Segmentation of Edge-On Galaxies

### v1

@@ Abstract
Edge-on galaxies have many important applications in galactic astrophysics, but they can be difficult to identify in vast amounts of astronomical data. To facilitate the search for them, we have developed a deep learning algorithm designed to identify and extract edge-on galaxies from astronomical images. We utilised a sample of edge-on spiral galaxies from the Galaxy Zoo database, retrieving the corresponding images from the Sloan Digital Sky Survey (SDSS). Our dataset comprises approximately $16,000$ galaxies, which we used to train the YOLOv5 algorithm for detection purposes. To isolate galaxies from their backgrounds, we trained the SCSS-Net neural network to generate segmentation masks. As a result, our algorithm detected $8,000$ edge-on galaxies, for which we compiled a catalogue including their parameters obtained from the SDSS database. We describe the basic properties of our sample, finding that most galaxies have redshifts $0.100.02<z<$ 0.10, have low values of $b/a$ and are mostly red, which is expected from edge-on galaxies and is consistent with our training sample, as well as other literature. The cutouts of the detected galaxies can be used for future studies and the algorithm can be applied to data from future surveys as well.
keywords:
galaxies: disc – galaxies: general – galaxies: fundamental parameters
pubyear: 2024pagerange: Advancing Galaxy Analysis: AI-Powered Detection and Segmentation of Edge-On GalaxiesAdvancing Galaxy Analysis: AI-Powered Detection and Segmentation of Edge-On Galaxies
@@ 1 Introduction
Edge-on galaxies are of great interest due to their unique orientation that enables us to study various phenomena. They come in many types and shapes, ranging from from bulge-dominated early-type galaxies to poorly understood bulgeless galaxies (Kautsch et al., 2006). Their structure is usually studied by decomposition into parts (e.g., Kormendy & Bruzual, 1978; Burstein, 1979; Hamabe et al., 1979; Tikhonov & Galazutdinova, 2005; Comerón et al., 2011; Salo et al., 2015), but they are also ideal for probing vertical and radial galactic structure (e.g., van der Kruit & Searle, 1981; Barteldrees & Dettmar, 1994; de Grijs & van der Kruit, 1996; Seth et al., 2005; Bizyaev et al., 2014), measuring the rotation curve (e.g., Makarov et al., 1999; Vergani et al., 2003; Kregel & van der Kruit, 2004; O’Brien et al., 2023) or investigating stellar haloes (e.g., Zibetti & Ferguson, 2004; Mouhcine et al., 2007; Mosenkov et al., 2020; Gilhuly et al., 2022). Edge-on galaxies are also suitable for observing peculiar features, such as galactic warps (e.g., Sancisi, 1976; Reshetnikov & Combes, 1998; Zee et al., 2022), disc truncation (e.g., Pohlen et al., 2000, 2002; van der Kruit, 2007; Martínez-Lombilla et al., 2019), galactic fountains (Bregman, 1980; Bregman & Houck, 1997; Fraternali et al., 2005; Hodges-Kluck & Bregman, 2012) or extraplanar gas (Pildis et al., 1994; Lee & Irwin, 1997; Levy et al., 2019; Reach et al., 2020) and extraplanar dust (e.g., Howk & Savage, 1997, 1999; Shinn, 2018; Mosenkov et al., 2022) among others.
With the increasing volume of data from various surveys, machine learning (ML) is becoming a necessity in galaxy analysis. There are many examples of using machine and deep learning for classifying galaxies based on their morphology (e.g., Storrie-Lombardi et al., 1992; Banerji et al., 2010; Dieleman et al., 2015; González et al., 2018; Barchi et al., 2020; Gharat & Dandawate, 2022, just to name a few), or from spectra, which enables the detection of active galactic nuclei (AGNs), star-forming or other galaxies (e.g., Shi et al., 2015; Zhang et al., 2019; Peruzzi et al., 2021; Wu et al., 2024). In the galactic context, ML can also be used to estimate redshift (e.g., Collister & Lahav, 2004; Bonnett, 2015; D’Isanto & Polsterer, 2018; Lin et al., 2022), identify galaxy-scale gravitational lensing (e.g., Cheng et al., 2020; Rezaei et al., 2022; Wilde et al., 2022; Andika et al., 2023), classify galaxy mergers (e.g., Pearson et al., 2019; Ferreira et al., 2020; Chang et al., 2022; Omori et al., 2023), estimate various galaxy parameters (e.g., Bonjean et al., 2019; Surana et al., 2020; Liew-Cain et al., 2021; Davidzon et al., 2022; Ghosh et al., 2022; Chu et al., 2024; La Torre et al., 2024), extract features (e.g., Abraham et al., 2018; Ghosh et al., 2020; Tarsitano et al., 2022; gord_2024) and for other applications.
@@
The application of ML algorithms specifically for detecting edge-on galaxies has just recently started to get explored. Savchenko et al. (2024) trained an artificial neural network (ANN) on edge-on galaxies extracted from the catalogue of genuine edge-on disc galaxies (EGIS, Bizyaev et al., 2014). Then they applied the algorithm on the data from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, Kaiser et al., 2010), where they detected $\sim 10^{5}$ galaxies with a detection rate of about $97\%$. Makarov et al. (2022) analysed the properties of this sample, finding that red sequence galaxies are thicker than the galaxies of the blue cloud and edge-on galaxies are systematically redder than the general population of galaxies seen at arbitrary angles. Usachev et al. (2024) applied this ANN algorithm on images from the Hubble Space Telescope COSMOS field (Scoville et al., 2007), where they detected 950 edge-on galaxies. They studied their scale-lengths and scale-heights, finding that their evolution is luminosity-dependent.
@@
In this work, we develop a deep learning algorithm which combines detection of edge-on, spiral galaxies followed by cutting them off (using segmentation) from the image automatically, so they are prepared for any further scientific analysis. This paper is structured as follows. In Section 2, we describe the selection of images, which were used for the algorithm training. In Section 3, we explain the algorithm and its performance in detail. In Section 4 we present our detected sample and its basic properties. In Section 5 we summarise our findings.
@@ 2 Data Selection
We are interested in detecting edge-on, spiral galaxies. Galaxies that are not completely edge-on, but are highly inclined are also appropriate, but for simplicity we will refer to our whole sample as edge-on. We selected our original sample from the Galaxy Zoo project (Lintott et al., 2008) and we downloaded the primary Galaxy Zoo 2 sample (Willett et al., 2013), which contains classifications of the $243,500$ galaxies in the main sample with spectroscopic redshifts. From this sample, following the criteria of Zee et al. (2022), we selected those galaxies, that at least one half of the volunteers voted to be edge-on and at least $80\%$ of the volunteers voted to be spiral galaxies. Then we retrieved the fits of these selected galaxies from the The Sloan Digital Sky Survey (SDSS) Data release 7 (Abazajian et al., 2009, DR7) Legacy Survey, which is the final data release of the SDSS Legacy Survey. We used all available fits files of all selected galaxies in all filters, so for most galaxies, we have several images. Those are treated as separate objects by the ML algorithm.
Once the images were selected ($15,500$ fits files), they were uploaded to the Zooniverse [^1] https://www.zooniverse.org/ platform, which became an important component of our data annotation process. The choice of Zooniverse was because of its proven track record in facilitating effective crowd-sourced scientific research, offering tools that simplify the complex tasks of data annotation for non-experts. This platform provided an intuitive interface for volunteers to interact with the data directly, ensuring that the annotations were accurate and verifiable. Volunteers were first presented with a brief tutorial explaining the characteristics of edge-on spiral galaxies and how to identify them. This tutorial included examples of images they would encounter and step-by-step instructions on how to mark the galaxies using the tools provided by the platform. The data collected through this method were then compiled and prepared for the next stages of our research, focusing on developing a machine learning algorithm to automate the detection of spiral galaxies. In the end, we have $16,701$ galaxies in $9,023$ fits files. An example of an annotated image is shown in Fig. 1.
3 Algorithm
Our algorithm for deep learning based extraction of edge-on galaxies, followed by acquisition of scientific data about detected objects, consists of three separated steps as part of main processing pipeline:
@@ 3 Algorithm
    Detection - in this step we acquire detection boxes of edge-on galaxies based on the results from YOLOv5 detection model, trained on selected data from SDSS survey.
@@
    Data Extraction - in this step, segmented masks are used to extract data from fits only for segmented edge-on galaxy, as well as cross-match detected galaxy with SDSS or other catalogues to extract selected parameters.
We provide more details on all three steps in next subsections.
3.1 YOLOv5 Training and Evaluation
@@
In our study, we utilised a dataset consisting of $9,023$ annotated images featuring galaxies. We divided this dataset into three subsets: training $(70\%)$, validation $(15\%)$, and testing $(15\%)$. The training set, comprising $6,314$ images, was used to train the YOLOv5 model (Jocher et al., 2020). The validation set, with $1,354$ images, helped evaluate the learning progress of the model. The testing set, which includes $1,353$ images, was used for the final model evaluation. It is important to note that if a galaxy appears in different filters, all versions of that galaxy are included in the same subset. This ensures that the same galaxy does not appear in multiple subsets. From a machine learning perspective, the different filters can be considered as data augmentations, which are very beneficial; however, this condition must be met to maintain the integrity of the data distribution.
@@
To prepare the data for YOLOv5, we focused on three key columns: metadata, subject_data, and annotations, extracting essential annotation information. A Python script generated the necessary text files for YOLOv5, containing labels and normalised coordinates $x$ and $y$, where the centre of the coordinate system $x,y=(0,0)$ is in the upper left corner of the image. The coordinates were normalised using the following formulas:
@@
The YOLOv5 detection model follows a structured architecture comprising three main parts: Backbone, Neck, and Head. The Backbone part (CSPDarkNet) extracts important features from images. The Neck part refines these features, often utilising Feature Pyramid Networks (FPN) to enhance spatial and semantic information. The Head part processes these outputs to produce bounding box predictions (detection boxes) and object classes. This architecture ensures efficient and accurate real-time object detection, often used in practice for image processing, making YOLOv5 a robust choice for our galaxy detection step (Jocher et al., 2020; Redmon et al., 2016; Vijayakumar & Vairavasundaram, 2024).
@@
  • Confusion Matrix: In general, it is a table or set of numbers which presents the performance of the prediction models by providing number of cases considered as True Positive (TP), False Positive (FP), False Negative (FN) and True Negative (TN) predictions. Due to the nature of detection tasks, where TNs represent the background without the expected detection and are thus not relevant, only the first three metrics are applicable and of interest, where:
@@
    $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (3)
@@
    $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (4)

### v2

# Deep Learning-Based Detection and Segmentation of Edge-On and Highly Inclined Galaxies

@@ Abstract
Edge-on galaxies have many important applications in galactic astrophysics, but they can be difficult to identify in vast amounts of astronomical data. To facilitate the search for them, we developed a deep learning algorithm designed to identify and extract edge-on galaxies from astronomical images. We utilised a sample of edge-on spiral galaxies from the Galaxy Zoo database, retrieving the corresponding images from the Sloan Digital Sky Survey (SDSS). Our dataset comprised $\sim 16,000$ galaxies, which we used to train the YOLOv5 algorithm for detection purposes. To isolate galaxies from their backgrounds, we trained the SCSS-Net neural network to generate segmentation masks. As a result, our algorithm detected $\sim 12,000$ edge-on galaxies with a high confidence, for which we compiled a catalog including their parameters obtained from the SDSS database. We described basic properties of our sample, finding that most galaxies have redshifts $0.100.02<z<$ 0.10, have low values of $b/a$ and are mostly red, which is expected from edge-on galaxies and is consistent with our training sample, as well as other literature. The cutouts of the detected galaxies can be used for future studies and the algorithm can be applied to data from future surveys as well.
facilities: Sloansoftware: APLpy (Robitaille & Bressert, 2012; Robitaille, 2019), astropy (Astropy Collaboration et al., 2013, 2018, 2022), astroquery (Ginsburg et al., 2019b), dustmaps (Green, 2018), ggpointdensity (https://github.com/LKremer/ggpointdensity), K-corrections calculator (Chilingarian et al., 2010; Chilingarian & Zolotukhin, 2012), matplotlib (Hunter, 2007), pandas (McKinney, 2010; The pandas development team, 2024), SciServer (Taghizadeh-Popp et al., 2020)
@@ 1 Introduction
Edge-on galaxies are of great interest due to their unique orientation that enables us to study various phenomena. Their high inclination allows for analysis in both radial and vertical direction, which was leveraged since early studies (Kormendy & Bruzual, 1978; Burstein, 1979; Hamabe et al., 1979; van der Kruit & Searle, 1981), usually with only a few well observed galaxies. They can be decomposed into parts, which enables analysis of different galactic components individually (e.g. Tikhonov & Galazutdinova, 2005; Mouhcine et al., 2007; Comerón et al., 2011; Gadotti & Sánchez-Janssen, 2012; Salo et al., 2015; Mosenkov et al., 2020; Gilhuly et al., 2022, and references therein). Edge-on galaxies can be of any type and shape, ranging from bulge-dominated early-type galaxies to poorly understood bulgeless galaxies (Kautsch et al., 2006), which makes them useful for studying various galactic components and peculiar features without biases due to inclination. Some features that are ideally observed in highly inclined galaxies include galactic warps, that were analysed by e.g. Sancisi (1976); Reshetnikov & Combes (1998); Zee et al. (2022), or galactic fountains, when disk gas is ejected into surrounding medium (Bregman, 1980; Bregman & Houck, 1997; Hodges-Kluck & Bregman, 2013; Putman et al., 2012). They are also ideal for observations of extraplanar gas (e.g. Reach et al., 2020, and references therein), extraplanar dust (e.g. Mosenkov et al., 2022, and references therein), polar bulges (Corsini et al., 2012; Reshetnikov et al., 2015; Mosenkov et al., 2024) and other features.
With the increasing volume of data from various surveys, machine learning (ML) is becoming a necessity in galaxy analysis. One of the most common applications of machine and deep learning is for classifying objects into categories. In the galactic context, different ML algorithms are used to classify galaxies based on their morphology into categories on the Hubble sequence (see Huertas-Company & Lanusse 2023 for a review) or based on their spectra, which enables categorization of galaxies into different types of active galactic nuclei (AGN), star-forming, composite and other types (e.g. Shi et al., 2015; Zhang et al., 2019; Peruzzi et al., 2021; Wu et al., 2024). Another type of ML application is training algorithms to predict various physical parameters of galaxies, such as photometric redshift (see Soo et al. 2023 for a review) or other physical properties (e.g. Bonjean et al., 2019; Surana et al., 2020; Liew-Cain et al., 2021; Davidzon et al., 2022; La Torre et al., 2024). ML is also efficient in detecting and extracting particular features of interest such as galactic bars (Abraham et al., 2018; Cavanagh & Bekki, 2020), galactic rings (Abraham et al., 2025), tidal features (e.g. Gordon et al., 2024, and references therein) among others.
@@
The application of ML algorithms specifically for detecting edge-on galaxies has just recently started to get explored. Savchenko et al. (2024) trained an artificial neural network (ANN) on edge-on galaxies extracted from the catalog of genuine edge-on disk galaxies (EGIS, Bizyaev et al., 2014). Then they applied the algorithm on the data from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, Kaiser et al., 2010), where they detected $\sim 10^{5}$ galaxies with a detection rate of about $97\%$. Makarov et al. (2022) analyzed the properties of this sample, finding that red sequence galaxies are thicker than the galaxies of the blue cloud and edge-on galaxies are systematically redder than the general population of galaxies seen at arbitrary angles. Usachev et al. (2024) applied this ANN algorithm on images from the Hubble Space Telescope COSMOS field (Scoville et al., 2007), where they detected 950 edge-on galaxies. They studied their scale-lengths and scale-heights, finding that their evolution is luminosity-dependent.
@@
The main motivation of this work is automatic detection and analysis of warped galaxies. In this paper, we present the first step, which is the detection and segmentation of edge-on galaxies, while in a future work we will focus on further training of the algorithm for warped galaxies. This paper is structured as follows. In Section 2, we describe the selection of images that were used for the algorithm training. In Section 3, we explain the processing workflow, algorithms, and its performance in detail. In Section 4 we present our detected sample and its basic properties. In Section 5 we summarize our findings.
@@ 2 Data Selection
We are interested in detecting edge-on, spiral galaxies. Galaxies that are not completely edge-on, but are highly inclined are also appropriate, but for simplicity we will refer to our whole sample as edge-on. We selected our original sample from the Galaxy Zoo project (Lintott et al., 2008) and we downloaded the primary Galaxy Zoo 2 sample (Willett et al., 2013), which contains classifications of the 243,500 galaxies in the main sample with spectroscopic redshifts. From this sample, following the criteria of Zee et al. (2022), we selected those galaxies, that at least one half of the volunteers voted to be edge-on and at least 80% of the volunteers voted to be spiral galaxies. We did not perform any additional visual inspection of the sample. Then we retrieved the fits of these selected galaxies from the The Sloan Digital Sky Survey (SDSS) Data release 7 (Abazajian et al., 2009, DR7) Legacy Survey, which is the final data release of the SDSS Legacy Survey. These FITS images have a size of 2048x1361 pixels, where 1 pixel = 0.396 arcsec. We downloaded them in all $ugriz$ bands, that have magnitude limits $u=22.0$ mag, $g=22.2$ mag, $r=22.2$ mag, $i=21.3$ mag, $z=20.5$ mag. We used all available fits files of all selected galaxies in all filters, so for most galaxies, we have several images. Those are treated as separate objects by the machine learning (ML) algorithm.
3 Methods
The general overview of our integrated approach is illustrated using the workflow diagram in Fig. 1, which depicts efficient processing of galaxy image data to ensure reliable preparation, modeling, and post-processing, developed in this work. This workflow is fundamental to understanding a systematic approach to the analysis of astronomical images and provides a clear path from raw data acquisition to the production of usable scientific data.
Figure 1: Diagram outlining the complete workflow used in the study, illustrating the process of data preparation, modeling and post-processing. Starting with the pre-processing phase, we transform the FITS data into JPEG format and annotate it, followed by splitting it into labeled datasets for machine learning training. The modeling phase involves training the YOLOv5 detection model and the SCSS-net segmentation model to identify and segment galaxies from astronomical images. The post-processing phase focuses on generating segmentation masks and extracting galaxy parameters for further analysis, demonstrating an integrated approach from raw data to usable scientific outputs.
The first part of our workflow under Preprocessing block defines the steps that were applied to provide input data for ML algorithms. Within data preparation, we prepared the images in suitable format, annotated them using the Zooniverse platform [^1] https://www.zooniverse.org/, and prepared training, validation, and testing sets for the start of the modeling phase.
The Modeling and Post-processing blocks can be recognized as the main algorithmic part of the processing pipeline and consists of three main procedural steps:
@@ 3 Methods
    Detection - in this step we acquire detection boxes of edge-on galaxies based on the results from the YOLOv5 detection model, trained on selected data from the SDSS survey.
@@
    Data Extraction - in this step, segmented masks are used to extract data from fits only for segmented edge-on galaxy, as well as cross-match detected galaxy with SDSS or other catalogs to extract selected parameters.
We provide more details on all steps of the workflow in the following subsections.
3.1 Preprocessing and Annotation of Selected Dataset
Once the images were selected ($15~{}500$ fits files), they were transformed to JPGs using the grayscale transformation of the FITSFigure wrapper for the transformation of FITS to JPG from APLpy package (Astronomical Plotting Library in Python) [^2] https://aplpy.readthedocs.io/en/stable/. In order for the images to be useful for annotation, we visualized them using APLpy’s function show_grayscale, applying a power stretch function with an exponent of $0.5$. The converted images have a size of 1000x727 pixels, with a horizontal and vertical resolution of 100 dpi, and bit depth 24. The images in JPG format were then uploaded to the Zooniverse platform, which became an important component of our data annotation process. The choice of Zooniverse was based on its proven track record of facilitating effective crowd-sourced scientific research, offering tools that simplify the complex tasks of data annotation for users. This platform provided an intuitive interface for volunteers to interact with the data directly, ensuring that the annotations were accurate and verifiable. Galaxy Zoo dataset is a good example of best practice for the use of the platform.
In our case, volunteers were first presented with a brief tutorial explaining the characteristics of edge-on spiral galaxies, as well as how to identify and annotate them. This tutorial included examples of images they would encounter and step-by-step instructions on how to mark the galaxies using the tools provided by the platform.
The project created in Zooniverse was available by public link, but in order to get annotations fast and reliably, we decided to go for annotation session with volunteers from students of Computer Science. We had 3 sessions with approximately 25 students (with 70 different volunteers in total), each 1.5 hours long. Together, the whole annotation process took 4.5 hours. Every session was in the same university classroom with computers and students also had the opportunity to consult any new specific case for annotation, so they were also able to learn during the process on how to proceed correctly with annotations.
Each image was annotated by one volunteer, so we did not need to take any additional steps to combine more bounding boxes for any image. However, it was possible that in the same image more than one edge-on galaxies occurred. Volunteers were instructed to provide a bounding box for each edge-on galaxy they were able to identify.
An example of an annotated image within the Zooniverse project interface is shown in Fig. 2.
Figure 2: An example of an interface with annotation of edge-on galaxies using a green bounding box by a volunteer, already added annotation of second galaxy in current image.
The data collected through the annotation step were then compiled and prepared for the next stages of our workflow, focusing on developing a machine learning algorithm to automate the detection of edge-on galaxies. In the end, we had $16~{}701$ galaxies identified in $9023$ fits files.
Before starting the detection model training, these files were randomly divided into three subsets: training set (70%), validation set (15%,) and testing set (15%).
3.2 Detection Task - Setup of YOLOv5 model
The prepared subsets were used for the training of the detection model, for which we decided to use the YOLO (You Only Look Once) detection network. It is a well-known one-shot (or single-stage) detector, which is used for detection of objects in images, originally developed for real-time monitoring cameras and fast classification of multiple objects in images, with no limit for the number of objects that can be found in one image. We have used the YOLOv5 version of this model (Jocher et al., 2020).
The YOLOv5 detection model follows a structured architecture comprising three main parts: Backbone, Neck, and Head. All details related to the particular implementation of YOLOv5, details on architecture and usage for the own custom datasets can be found in its documentation (see Architecture Summary part in documentation guidelines of YOLOv5) [^3] https://docs.ultralytics.com/yolov5.
The Backbone part extracts features from input images and learns relevant spatial and semantic information using CSPDarknet53, an enhanced version of Darknet53 convolutional network that incorporates Cross Stage Partial (CSP) connections. At the start, YOLOv5 introduces a Focus layer, which reorganizes pixel values to reduce spatial dimensions while preserving important details. In the end of the Backbone part, features of different types and scales are efficiently extracted. The Neck provides a fusion of extracted features, ensuring that information from different scales is effectively combined. YOLOv5 integrates a Path Aggregation Network (PANet), which enhances the flow of information between different feature levels, including the application of the Feature Pyramid Network (FPN) to strengthen multi-scale feature representation. This combination allows the model to detect objects of varying sizes with improved accuracy. The Head part processes these outputs to produce detections (bounding boxes). This architecture ensures efficient and accurate real-time object detection, often used in practice for image processing, making YOLOv5 a robust choice for our galaxy detection step (Jocher et al., 2020; Redmon et al., 2016; Vijayakumar & Vairavasundaram, 2024).
@@ 3.2 Detection Task - Setup of YOLOv5 model
For our detection training, 6314 images were used to train the model. The validation set with 1354 images helped to evaluate the learning progress of the model during training and optimization of the model. The test set, which includes 1353 images, was used for the final evaluation of the model. It is important to note that if a galaxy appears in different filters, all versions of that galaxy are included in the same subset. This ensures that the same galaxy does not appear in multiple subsets. This condition must be met to maintain the integrity of the data distribution. From a machine learning perspective, the different filters within one subset can be considered as data augmentations, which are very beneficial for including variations in target images for future predictions on unseen data.
@@
To prepare the data for YOLOv5, we focused on three key columns: metadata, subject_data, and annotations, extracting essential annotation information. A Python script generated the necessary text files for YOLOv5, containing labels and normalized coordinates $x$ and $y$, where the center of the coordinate system $x,y=(0,0)$ is in the upper left corner of the image. The coordinates were normalized using the following formulas:
@@
  • Confusion Matrix: In general, it is a table or set of numbers which presents the performance of the prediction models by providing number of cases considered as True Positive (TP), False Positive (FP), False Negative (FN) and True Negative (TN) predictions. Due to characteristic and output of detection tasks (TN is considered as background of image without expected detection and is out of scope for such a task), only the first three are applicable and interesting for detection, where:
@@
[…]
@@
[…]
@@
3.3 Segmentation Task Based on U-Net Architecture
Our main motivation behind this work is the analysis of warps in edge-on galaxies, which we plan to present in future work. Therefore, to analyze our detected galaxies, we need to segment the image and extract the galaxies while removing nearby objects that might contaminate them. There are several methods of image segmentation, such as thresholding, when the object is selected based on having a significantly higher pixel intensity value than the background (e.g., Lotz et al., 2004; Zee et al., 2022), using isophotes to define the extent of the galaxy (e.g., Hao et al., 2006; Reshetnikov et al., 2016), deep learning (e.g., Boucaud et al., 2020; Farias et al., 2020) and other methods (for a comprehensive list of methods and their comparisons see Xu & Zhu, 2024). Each method has its advantages and issues, but in all cases determining exactly the threshold where the galaxy ends and the background starts can be challenging.
We decided to extract the galaxies using our own segmentation algorithm. We used the bounding-box coordinates obtained from the YOLOv5 detection to crop galaxies with a close background from the original images. This procedure ensured that the segmented images contained only relevant galaxy data, thus optimizing the segmentation process.
The process of creating high-quality segmentation masks for training the model from cropped galaxy images was time-consuming, mainly due to the manual adjustment of parameters. Each iteration involved careful parameter settings for thresholding and morphological operations such as erosion and dilation. After setting these parameters, we manually checked the resulting masks to ensure that they met our standards for clarity and definition of galaxy structures.
To automate the generation of final segmentation masks, we used SCSS-Net, a convolutional neural network inspired by the U-Net architecture, originally designed to segment solar corona structures (Mackovjak et al., 2021). This versatile model was adapted for our specific requirement and re-trained for segmentation of edge-on galaxies.
U-Net is a highly efficient convolutional neural network originally developed for medical image segmentation. Its architecture, designed to work effectively even with a limited amount of data, is characterized by its distinctive U-shape, comprising two main components: an encoder and a decoder (Ronneberger et al., 2015).
Encoder: The encoder is responsible for progressively capturing the context of the image, which involves reducing spatial dimensions while increasing depth. This process helps to extract and compress feature information, which is crucial for understanding the complex structures within the image.
Decoder: In contrast, the decoder part of the network works by upscaling the feature map to reconstruct the image output. Eventually, it recovers the spatial dimensions and detail of the image, ensuring that the important features extracted by the encoder are used to generate a precise and detailed segmentation. In our case, the input is the original galaxy image with our manually created mask and the output is the mask created by the algorithm.
A more detailed description of our version of the U-Net model (SCSS-Net) used for edge-on galaxy segmentation is shown in Figure 4. This architecture not only allows for detailed image segmentation, but also ensures that the network learns from a relatively small amount of data, making it particularly powerful in fields where high-quality data is scarce.
Figure 4: The architecture of segmentation model, originally designed for solar corona structures segmentation as SCSS-Net, is inspired by U-Net configuration. For our segmentation of edge-on galaxies, input image (for best model 64x64) of bounding box with edge-on galaxy is at the start of this Encoder-Decoder architecture. The encoder (left part of architecture) is composed of five convolution blocks connected using max pooling layers (with pool size of 2x2 pixels). Convolutional blocks contain two convolutional layers, batch normalization and ReLU activation functions (see ConvBlock internal structure depiction on the right side of the figure). The convolutional layers in blocks double their number of filters for each depth level - from 32 at the start to 512 at the bottom of the encoder part. The decoder (right part of the architecture) contains four deconvolutional blocks connected using up-sampling with a pool size of 2×2 pixels. Deconvolution blocks (see DeconvBlock internal structure depicted on the right side of the figure) consist of two transpose convolution layers, batch normalizations, and the ReLU activation functions. Deconvolutional layers progressively halving number of filters from 256 to 32. The deeper layers also contain dropout regularization layers. Additionaly, concatenation layers are used to concatenate feature maps within the same level of scale (granularity), which helps to give the localization information from encoder to decoder. The output layer convolutional layer and sigmoid activation function, providing segmentation output mask for the same size as input image.
For the learning of the segmentation model, we again divided our data set into training, validation, and test sets, and tuned basic hyperparameters such as image size, batch size, and number of epochs. Due to our previous experience with SCSS-Net, we knew that for this type of task approximately 1000 training images would be sufficient for good segmentation results. Therefore, we prepared enough manually created masks for this purpose. After the preparation of masks, the input dataset with 1494 masks of galaxy images was divided into three subsets: 1091 were used as a training set, 273 masks as validation set, and 130 masks as test set. Due to the small size of our images ($64\times 64$ pixels), the initial training phase was efficient. The whole model was trained from scratch. We tested different image sizes for this purpose (including larger bounding-box inputs, e.g. $128\times 128$ or $256\times 256$), but the best results were achieved using $64\times 64$. We aimed for 1000 epochs with a batch size of 20, but concluded training at epoch 963 as performance gains plateaued.
  • Intersection over Union (IoU): Also known as the Jaccard Index, IoU is a statistical tool used to quantify the percent overlap between the target mask and the prediction output of the model. It is calculated as the area of overlap between the predicted segmentation and the ground truth divided by the area of union between the predicted and ground truth. This metric is particularly useful for determining how well the predicted boundaries align with the true boundaries.
[…]
3.4 Evaluation of YOLO-based Detection Results